OpenStudy (apricot):

Which of the following is equivalent to ½(4x - 6y) ? 8x + 12y (½) 4x - 6y 2x - 3y 2x + 3y

Vocaloid (vocaloid):

hint: use the distributive property

Vocaloid (vocaloid):

A(B-C) = A*B - A*C ½(4x - 6y) = ________ fill in the blank

Vocaloid (vocaloid):

still there?

OpenStudy (apricot):

½ (4x) - ½ (6y)?

Vocaloid (vocaloid):

good, you're almost there 1/2 * 4x = ?

OpenStudy (apricot):

is it 2x - 3y? I originally thought this but I wasn't sure

Vocaloid (vocaloid):

yup that scorrect!

OpenStudy (apricot):

thank you! :)